An executive, who just turned forty five years of age, is considering two jobs offers. The first would pay $250,000 at the end of next three years. The second would pay $100,000 at the end of next five years plus retirement account would be set up that would pay a bonus of $3,000,000 on the executive’s sixty-fifth birthday. The interest rate is 12 percent. If money is only considering which offer should be accepted?
1
Expert's answer
2015-04-25T11:00:31-0400
i = 12%
If the first would pay $250,000 at the end of next three years, then on the executive’s sixty-fifth birthday he will have FV1 = (250,000*1.12 + 250,000*1.12^2 + 250,000*1.12^3)*1.12^17 = $6,487,255.14. If the second would pay $100,000 at the end of next five years plus retirement account would be set up that would pay a bonus of $3,000,000 on the executive’s sixty-fifth birthday, then FV2 = (100,000*1.12 + 100,000*1.12^2 + 100,000*1.12^3 + 100,000*1.12^4 + 100,000*1.12^5)*1.12^15 + 3,000,000 = $6,894,545.51.
So, the second offer should be accepted.